# Next-to-leading-order electroweak correction to $H\to Z^{0}\gamma$

### v1

@@ Abstract
Summary. In this work we have calculated the NLO electroweak correction to the rare decay process $H\to Z^{0}\gamma$ within the on-shell renormalization scheme. To assess the theoretical uncertainty, we present the predictions of the partial width and the branching fraction using several different $\alpha$ schemes. In contrast to the tiny NLO QCD correction, the magnitude of the NLO electroweak correction turns out to become sizable, reaching $7\%$ of the LO result in both the $\alpha(0)$ and $\alpha(M_{Z})$ schemes. After including the ${\cal O}(\alpha)$ correction, the predictions from various $\alpha$ schemes tend to converge to each other, which indicates that incorporating the NLO electroweak correction has significantly stabilized the SM prediction. Our most accurate prediction is then ${\cal B}[H\to Z^{0}\gamma]=(1.55\pm 0.06)\times 10^{-3}$, with the uncertainty predominantly stemming from the error of the full width of Higgs. The relative error from varying the $\alpha$ schemes is less than $2\%$. Although this most accurate SM prediction is significantly lower than the measured value, it might be premature to claim that one has to invoke the new physics to resolve this discrepancy. Likely the problem mainly lies on the experimental side. Improved measurement of this rare decay process at HL-LHC, and independent measurements at the prospective Higgs factories such as the CEPC and the FCC, are crucial.
@@ Acknowledgements.
We thank Yingsheng Huang for suggesting us to consider this project. We are also indebted to Wen Chen and Yingsheng Huang for discussions. Feynman diagrams in this work are drawn with the aid of JaxoDraw Binosi:2008ig . The work of W.-L. S. is supported by the NNSFC under Grant No. 12375079, and the Natural Science Foundation of ChongQing under Grant No. CSTB2023 NSCQ-MSX0132. The work of F. F. is supported by the NNSFC Grant No. 12275353. The work of Y. J. is supported in part by the NNSFC Grants No. 11925506.
Note added. While we were finalizing this work, a preprint by Chen, Chen, Qiao and Zhu has recently appeared Chen:2024vyn . Our results for the $\alpha({M_{Z}})$ and $G_{\mu}$ schemes are compatible with their mixed 1 and mixed 2 schemes, provided that the same input parameters are used. However, our result for the $\alpha(0)$ scheme slightly differs from theirs, probably due to the different treatment of the light quark contribution to $Z_{e}$.
@@ Appendix A The expressions of $\widetilde{T}_{5}$ from four different $\alpha$ schemes
As aforementioned, there are some freedoms in the on-shell renormalization scheme to handle the charge renormalization. In this appendix, we present the expressions of the form factor $\widetilde{T}_{5}$ in four different schemes, which differ in choosing the values of the QED coupling constant. Throughout this work we always retain the QED coupling associated with the photon emission vertex to be $\alpha(0)$. Plugging these equations into (3), we then obtain the predicted partial widths associated in each scheme, as enumerated in Table 1.
@@ A.1 $\alpha(0)$ scheme
$\widetilde{T}_{5}=1.546\,\alpha^{3/2}(0)\bigg{[}1+\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})+4.070\;\frac{\alpha(0)}{\pi}$ (8)
@@
which depends on the light charged leptons but not on light quarks (whose effect has been included in $\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})$). As can be clearly seen, the prefactor accompanying $\alpha(0)$ is about two orders-of-magnitude greater than that accompanying $\alpha_{s}$, which explains why the electroweak correction is much more important than the QCD correction.
@@ A.2 $\alpha(M_{Z})$ scheme
In this scheme, we choose the QED coupling constants associated with two vertices other than the photon emission vertex in Fig. 1 to be $\alpha(M_{Z})$. Making the following replacement in (8):
$\alpha(0)$$\to$ $\alpha(M_{Z})\bigg{[}1-\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})-\Delta\alpha_{\ell}(M_{Z})\bigg{]}$ (9)
$=$ $\alpha(M_{Z})\bigg{[}1-0.02764-18.132\,\frac{\alpha(M_{Z})}{\pi}\bigg{]},\phantom{xx}$
where $\Delta\alpha_{\ell}(M_{Z})=\Pi^{\gamma\gamma}_{f=e,\mu,\tau}(0)-{\rm Re}\,\Pi^{\gamma\gamma}_{f=e,\mu,\tau}(M_{Z})$, we then obtain
@@
Note $\widetilde{T}_{5}$ is insensitive of the mass of all light fermions.
@@ A.3 $G_{\mu}$ scheme
$\alpha(0)$$\to$ $\alpha_{G_{\mu}}\bigg{[}1-\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})+\Delta\alpha_{lq}(M_{Z})-\Delta r\bigg{]}$ (11)
$=$ $\alpha_{G_{\mu}}\bigg{[}1-0.02764-5.635\,\frac{\alpha_{G_{\mu}}}{\pi}\bigg{]},\phantom{xx}$
with $\Delta\alpha_{lq}(M_{Z})=\Pi^{\gamma\gamma}_{f=u,d,s,c,b}(0)-{\rm Re}\,\Pi^{\gamma\gamma}_{f=u,d,s,c,b}(M_{Z})$, we obtain
@@
As in $\alpha(M_{Z})$ scheme, $\widetilde{T}_{5}$ in $G_{\mu}$ scheme is also independent of the masses of all light fermions.

### v3

@@ Abstract
Summary. In this work we have calculated the NLO electroweak correction to the rare decay process $H\to Z^{0}\gamma$ within the on-shell renormalization scheme. To assess the theoretical uncertainty, we present the predictions of the partial width and the branching fraction using several different $\alpha$ schemes. In contrast to the tiny NLO QCD correction, the magnitude of the NLO electroweak correction turns out to become sizable, reaching $7\%$ of the LO result in the $\alpha(M_{Z})$ schemes. After including the ${\cal O}(\alpha)$ correction, the predictions from various $\alpha$ schemes tend to converge to each other, which indicates that incorporating the NLO electroweak correction has significantly stabilized the SM prediction. Our most accurate prediction is then ${\cal B}[H\to Z^{0}\gamma]=(1.55\pm 0.06)\times 10^{-3}$, with the uncertainty predominantly stemming from the error of the full width of Higgs. The relative error from varying the $\alpha$ schemes is less than $2\%$. Although this state-of-the-art SM prediction is significantly lower than the measured value, it is still premature to claim that one has to invoke some BSM scenarios to resolve this discrepancy. Likely the problem mainly lies on the experimental side. Improved measurements of this rare decay process at HL-LHC, and independent measurements at the prospective Higgs factories such as CEPC and FCC-ee, will be crucial.
@@ Acknowledgements.
We thank Yingsheng Huang for suggesting us to consider this project. We are also indebted to Wen Chen and Yingsheng Huang for discussions. Feynman diagrams in this work are drawn with the aid of JaxoDraw Binosi:2008ig . The work of W.-L. S. is supported by the NNSFC under Grant No. 12375079, and the Natural Science Foundation of ChongQing under Grant No. CSTB2023 NSCQ-MSX0132. The work of F. F. is supported by the NNSFC under Grant No. 12275353. The work of Y. J. is supported in part by the NNSFC under Grant No. 11925506.
Note added. While we were finalizing this work, a preprint by Chen, Chen, Qiao and Zhu has recently appeared Chen:2024vyn . Our results in $\alpha({M_{Z}})$ and $G_{\mu}$ schemes are compatible with their mixed 1 and mixed 2 schemes, provided that the same input parameters are used. However, our result in $\alpha(0)$ scheme slightly differs from theirs, probably due to the different treatment of the light quark contribution to $Z_{e}$.
@@ Appendix A The expressions of $\widetilde{T}_{5}$ from four different $\alpha$ schemes
Throughout this work we always retain the QED coupling associated with the photon emission vertex to be $\alpha(0)$. As aforementioned, there still exists some freedom in the on-shell renormalization scheme to handle the charge renormalization. In this appendix, we present the expressions of the form factor $\widetilde{T}_{5}$ in four different $\alpha$ schemes, which differ in choosing the values of the QED coupling constant in other vertices. Plugging these equations into (3), we then obtain the predicted partial widths associated with each scheme, as enumerated in Table 1.
@@ A.1 $\alpha(0)$ scheme
$\widetilde{T}_{5}=1.546\,\alpha^{3/2}(0)\bigg{[}1+\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})-0.535\;\frac{\alpha(0)}{\pi}$ (9)
@@
which depends on both charged leptons and light quarks, with the effects of the latter encapsulated in $\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})$. As can be clearly seen, the prefactor accompanying $\alpha(0)$ is about two orders-of-magnitude greater than that accompanying $\alpha_{s}$, which explains why the electroweak correction is much more important than the QCD correction.
@@ A.2 $\alpha(M_{Z})$ scheme
In this scheme, we choose the QED coupling constants associated with two vertices other than the photon emission vertex in Fig. 1 to be $\alpha(M_{Z})$. With the aid of (6) and (8a), making the following substitution in (9):
$\alpha(0)$$\to$ $\alpha(M_{Z})\bigg{[}1-\Delta{\alpha}(M_{Z})\bigg{]}$ (10)
$\to$ $\alpha(M_{Z})\bigg{[}1-\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})-\Delta\alpha_{\ell}(M_{Z})\bigg{]}$
$=$ $\alpha(M_{Z})\bigg{[}1-0.02764-13.527\,\frac{\alpha(M_{Z})}{\pi}\bigg{]},\phantom{xx}$
we then obtain
@@
To be compatible with the level of accuracy in the computation of $\delta{Z_{e}}\big{|}_{\alpha(0)}$ in $\alpha(0)$ scheme, we have utilized the one-loop expression of $\Delta\alpha_{\ell}(M_{Z})$ in (10), which reads
$\Delta\alpha_{\ell}(M_{Z})=\frac{\alpha(0)}{3\pi}\bigg{(}\ln\frac{M_{Z}^{2}}{m_{e}^{2}}+\ln\frac{M_{Z}^{2}}{m_{\mu}^{2}}+\ln\frac{M_{Z}^{2}}{m_{\tau}^{2}}-5\bigg{)}.$ (12)
It might be worth emphasizing that $\widetilde{T}_{5}$ in (11) is insensitive to the masses of all light quarks and charged leptons.
@@ A.3 $G_{\mu}$ scheme
$\alpha(0)$$\to$ $\alpha_{G_{\mu}}\bigg{[}1-\Delta{\alpha_{\rm had}^{(5)}}(M_{Z})+\Delta\alpha_{q}(M_{Z})-\Delta r\bigg{]}$ (13)
$=$ $\alpha_{G_{\mu}}\bigg{[}1-0.02764-1.030\,\frac{\alpha_{G_{\mu}}}{\pi}\bigg{]},\phantom{xx}$
we then obtain
@@
Note that in (13) we have utilized the one-loop values for $\Delta\alpha_{q}(M_{Z})$ and $\Delta r$. The explicit expression of $\Delta r$ is too complicated to present here, which can instead be found in Eq. (8.14) of Denner:1991kt . It is worth emphasizing that since both $\Delta\alpha_{q}(M_{Z})$ and $\Delta r$ possess identical sensitivity to the light quark masses, their difference in (13) is not sensitive to nonperturbative hadronic effect, thus amenable to a perturbative treatment. As in $\alpha(M_{Z})$ scheme, $\widetilde{T}_{5}$ in $G_{\mu}$ scheme is also insensitive to the masses of light quarks and charged fermions.